determine the exactness and obtain the general solutions.

1. (π‘₯ + 𝑦2 )𝑑π‘₯ + (2π‘₯𝑦 + 𝑦2)𝑑𝑦 = 0

2. (2π‘₯𝑦3 + π‘₯2 )𝑑π‘₯ + (3π‘₯2𝑦2 + 𝑦)𝑑𝑦 = 0

3. (𝑒2y + π‘₯)𝑑π‘₯ + (2π‘₯𝑒2y + 1)𝑑𝑦 = 0Β 


1
Expert's answer
2021-10-06T00:24:36-0400

Let us determine the exactness and obtain the general solutions.


1. Consider the equation "(\ud835\udc65 + \ud835\udc66^2 )\ud835\udc51\ud835\udc65 + (2\ud835\udc65\ud835\udc66 + \ud835\udc66^2)\ud835\udc51\ud835\udc66 = 0". Since "\frac{{\partial (x+y^2)}}{\partial y}=2y=\frac{{\partial (2xy+y^2)}}{\partial x}," we conclude that the differential equation is exact. Then "\frac{{\partial U}}{\partial x}=x+y^2, \frac{{\partial U}}{\partial y}=2xy+y^2." It follows that "U=\frac{x^2}2+y^2x+C(y)," and hence "\frac{{\partial U}}{\partial y}=2yx+C'(y)=2xy+y^2." Therefore, "C'(y)=y^2," and thus "C(y)=\frac{y^3}3+C." We conclude that the general solution of the differential equation is "\frac{x^2}2+y^2x+\frac{y^3}3=C."


2. Consider the equation "(2\ud835\udc65\ud835\udc66^3 + \ud835\udc65^2 )\ud835\udc51\ud835\udc65 + (3\ud835\udc65^2\ud835\udc66^2 + \ud835\udc66)\ud835\udc51\ud835\udc66 = 0". Since "\frac{{\partial (2𝑥𝑦^3 + 𝑥^2)}}{\partial y}=6xy^2=\frac{{\partial (3𝑥^2𝑦^2 + 𝑦)}}{\partial x}," we conclude that the differential equation is exact. Then "\frac{{\partial U}}{\partial x}=2𝑥𝑦^3 + 𝑥^2, \frac{{\partial U}}{\partial y}=3𝑥^2𝑦^2 + 𝑦." It follows that "U=x^2y^3+\frac{x^3}3+C(y)," and hence "\frac{{\partial U}}{\partial y}=3x^2y^2+C'(y)=3x^2y^2+y." Therefore, "C'(y)=y," and thus "C(y)=\frac{y^2}2+C." We conclude that the general solution of the differential equation is "x^2y^3+\frac{x^3}3+\frac{y^2}2=C."


3. Consider the equation "(\ud835\udc52^{2y} + \ud835\udc65)\ud835\udc51\ud835\udc65 + (2\ud835\udc65\ud835\udc52^{2y} + 1)\ud835\udc51\ud835\udc66 = 0". Since "\frac{{\partial (𝑒^{2y} + 𝑥)}}{\partial y}=2e^{2y}=\frac{{\partial (2𝑥𝑒^{2y} + 1)}}{\partial x}," we conclude that the differential equation is exact. Then "\frac{{\partial U}}{\partial x}=𝑒^{2y} + 𝑥, \frac{{\partial U}}{\partial y}=2𝑥𝑒^{2y} + 1." It follows that "U=xe^{2y}+\frac{x^2}2+C(y)," and hence "\frac{{\partial U}}{\partial y}=2xe^{2y}+C'(y)=2xe^{2y}+1." Therefore, "C'(y)=1," and thus "C(y)=y+C." We conclude that the general solution of the differential equation is "xe^{2y}+\frac{x^2}2+y=C."